OpenStudy (anonymous):

Write a polynomial function of minimum degree with real coefficients whose zeros include those listed. Write the polynomial in standard form. 5, -3, and -1 + 3i

OpenStudy (anonymous):

start with \[(x-5)(x+3)(x-(-1+3i))(x-(-1-3i))\] and multiply out that is the hard way to do it, but there is a somewhat easier way

OpenStudy (anonymous):

whats easier?

OpenStudy (anonymous):

you have no choice but to multiply the first product is not hard \[(x-5)(x+3)=x^2-2x-15\] it is the second product that has an easier way to find then straight out multiplication

OpenStudy (anonymous):

yeah that's the part i dont understand

OpenStudy (anonymous):

you know one of the zeros is \(-1+3i\) so you can find the quadratic polynomial using one of two simple methods

OpenStudy (anonymous):

start with \[x=-1+3i\] and work backwards

OpenStudy (anonymous):

\[x=-1+3i\] \[x+1=3i\] \[(x+1)^3=(3i)^2=-9\] \[x^2+2x+1=-9\] \[x^2+2x+10=0\]

OpenStudy (anonymous):

that is the polynomial with zeros \(-1+3i\) and \(-1+3i\)

OpenStudy (anonymous):

alternatively, you can simply remember that if \(a+bi\) is the zero of a quadratic polynomial with leading coefficient 1, that polynomial is \[x^2-2a+(a^2+b^2)\]

OpenStudy (anonymous):

so then just factor the two polynomials?

OpenStudy (anonymous):

in your case \(a=-1,b=3\) so the polynomial must be \[x^2+2x+10\]

OpenStudy (anonymous):

no no don't "factor" you have it in factored form to begin with your job is to multiply out, not factor

OpenStudy (anonymous):

\[(x^2-2x-15)(x^2+2x+10)\] is what you have to compute to write the polynomial in standard form

OpenStudy (anonymous):

thats what i meant! lol sorry